# Benchmarking Quantum Reinforcement Learning

### v1

@@ 1 Introduction
In the search for heuristic quantum advantage in QRL, we encounter issues similar to classical RL: sensitivity to hyperparameter choices, various randomness sources, and even random seed and codebase (Henderson et al., 2018; Jordan et al., 2024a). QRL faces similar, if not more severe, reproducibility issues than traditional RL (Bowles et al., 2024; Franz et al., 2023) due to additional randomness. Therefore, any claim of QRL’s superiority over classical RL should be taken with great care.
@@
In addition to common sources of randomness known in RL (Henderson et al., 2018), such as random weight initialization, randomness in the environment, etc., QRL is subject to additional sources of randomness such as shot noise or hardware imperfections due to current limitations. Therefore, performance comparisons need to be based on a sound statistical evaluation. However, most studies perform only a small, potentially insufficient number of training runs for robust inferences under stochasticity. As a result, statements are potentially misleading or insignificant. Fig. 2 exemplifies the comparison of two RL algorithms w.r.t some threshold on the evaluated return. It illustrates common flaws prevalent in the QRL literature, e.g., averaging learning curves over only 5 seeds, inconsistency in statistical ranges, etc. Fig. 2 indicates that algorithm 1 is more sample-efficient. However, the interquartile ranges (shaded areas) estimated with a much larger sample size of $100$ runs rather support the opposite statement.
@@
The remainder of this paper is structured as follows. Sec. 2 reviews related work. Sec. 3 provides background on (Quantum) RL. Next, Sec. 4 introduces our statistical estimator. Sec. 5 describes the experimental setup and introduces a novel environment with flexible complexity. Sec. 6 evaluates QRL using our statistical estimator. We discuss the implications on quantum advantage in Section 7.
@@ 2 Related Work
For classical deep RL, in general, no rigorous analysis of algorithms is possible, necessitating a computational approach to algorithmic comparison and representative benchmarking environments (Brockman et al., 2016; Todorov et al., 2012; Tassa et al., 2018a) for mature implementations of widely-used deep RL algorithms (Hill et al., 2018; Duan et al., 2016; Fujimoto et al., 2019; Wang et al., 2019b). However, benchmarking results are heavily influenced by (sometimes seemingly trivial) implementation details (Engstrom et al., 2019; Andrychowicz et al., 2021; Huang et al., 2022), as well as randomness in environment transitions and network initialization (Henderson et al., 2018). Research has focused on determining the necessary number of seeds for robust comparison (Agarwal et al., 2021; Colas et al., 2018) and on adopting statistical analysis over simple point estimates (Agarwal et al., 2021; Colas et al., 2018; Patterson et al., 2024) (e.g., reporting the interquartile mean instead of the average performance). However, benchmarking algorithms with increased number of seeds (Laskin et al., 2019; Gorsane et al., 2022; Bettini et al., 2024) can quickly become a bottleneck (Jordan et al., 2024b), especially for computationally-intense algorithms, like model-based (Wang et al., 2019b) or safe RL (Zhao et al., 2023).
@@
QRL can solve artificial tasks (without practical relevance) with exponentially smaller sample complexity compared to any known classical algorithm. Recently, (Jerbi et al., 2021; Liu et al., 2021) constructed artificial problems which are widely believed to be classically intractable. When the type of the algorithm is fixed to e.g. policy iteration, quantum versions exist with polynomially reduced sample complexity (Wang et al., 2021; Ganguly et al., 2023; Zhong et al., 2023). However, these algorithms require resources that far exceed the capabilities of current quantum hardware. An alternative line of research (Chen et al., 2020; Skolik et al., 2022; Lockwood & Si, 2020), more aligned with the limitations of current hardware, substitutes the classical neural network used as a function approximator for policy and value function in classical RL algorithms by a variational quantum circuit VQC (Bharti et al., 2022). Here, Jerbi et al. (2021) construct artificial problems based on the same VQC architecture later used by the learner. The superiority found empirically for the QRL algorithm can then be attributed to the inductive bias introduced in the problem. Several studies have compared classical RL and QRL on toy problems inspired by real-world tasks. Some of these studies (Chen et al., 2020; Drǎgan et al., 2024; Reers, 2023; Hohenfeld et al., 2024) have documented empirical superiority of QRL on metrics closely related to sample complexity. This work proposes a robust methodology to compare these types of algorithms.
@@
We compare two function approximation approaches in these algorithms: (i) the standard version with a classical DNN and (ii) the quantum version, which has a variational quantum circuit between two small classical fully-connected layers. This structure is called hybrid classical-quantum, as depicted in Figure 3. As discussed in Section D.2, the classical layers have few trainable parameters, and small fully classical DNNs are not competitive, so the hybrid models’ performance originates from the quantum sub-module. The VQC acts like a quantum counterpart to a classical DNN, with trainable weights parameterizing unitary operations, usually single-qubit rotations (Bharti et al., 2022). The quantum state is measured to approximate complex functions. Past research suggests VQCs may have advantages over DNNs, including better accuracy for certain tasks (Liu et al., 2021) and smaller model size (Chen et al., 2020). More details on VQCs and the ansatz used here are shown in Section D.1. Such models have been used in value-based (Chen et al., 2020; Skolik et al., 2022) and policy-based (Jerbi et al., 2021; Meyer et al., 2023) RL routines.
@@ 5 Experimental Setup
As QRL algorithms exhibit some robustness against hardware noise (Skolik et al., 2023) (which is expected to further decrease in the future (Kim et al., 2023; Acharya et al., 2024)) we decided to perform all experiments in a noise-free simulation. We reduced the impact of parameter initialization by considering $100$ random seeds per environment and model configuration. We initialized the classical neural networks using He initialization and the VQC parameters uniformly at random within $\left[0,2\pi\right]$. We tuned the hyperparameters, see Appendix C, before running the actual experiments. The configurations of the models are reported in terms of three metrics: (i) model width, i.e., the number of neurons in hidden layers for classical DNNs, and in case of quantum architectures the number of qubits; (ii) model depth, i.e., the number of hidden layers for classical networks, and the number of layers of the quantum models; and (iii) model complexity, parameterized by the number of trainable parameters, which will be justified further below. The reported configurations were chosen by an extensive ablation study, see Appendix D.
@@ 6 Experiments
We evaluate different (Q)RL algorithms on our novel BeamManagement6G environment using our proposed statistical sample complexity estimator in Sec. 6.1. Sec. 6.2 studies the relationship between sample complexity and (RL) model complexity. In Sec. 6.3 we compare the performance of the trained policies.
@@ 6.2 Scaling with Model Complexity
We now explore the relationship between sample efficiency and the complexity of the RL and QRL model, respectively. There exist different measures for model complexity both in the classical (Hu et al., 2021) and quantum (Abbas et al., 2021) domain. In this work we define model complexity based on the number of trainable parameters. This one-dimension projection in the model space provides a sufficiently reliable metric for the arguments below.
@@
At this point, two questions arise which are addressed in the following: (i) Are models with lower complexity but similar performance generally preferrable, i.e., how should sample and model complexity be balanced? (ii) Is outperfomance of the quantum model achievable by further increasing qubit numbers? The first question (i) cannot be answered in general as it depends on concrete practical considerations and objectives. If the primary goal is to minimize the number of agent-environment interactions, regardless of training and inference costs, the classical models are superior. However, when memory requirements are considered, the reduced parameter count of the quantum-enhanced model might be advantageous. Assuming scalability of this approach, also runtime improvements are conceivable. The second question (ii) cannot be answered conclusively at the current stage. The saturation behavior in Figure 7 suggests that further scaling the quantum model could lead to comparable or superior performance relative to the largest classical model. We stress that a naive extrapolation to higher qubit numbers might be limited in practice by trainability issues (Larocca et al., 2024). Moreover, it is computationally prohibitive to generate simulation results beyond $14$ qubits due to the substantial overhead required for statistically robust results as summarized in Table 3 in Section D.2.
@@ 6.3 Performance of Trained Policies
As a consequence of different trajectories, the maximum achievable intensity varies. Both RL strategies produce similar-looking histograms of the received intensities. The reported mean values of the intensity of the quantum model (green) and the classical model (blue) do not indicate significantly different behavior by visual inspection. Furthermore, we observed comparable results for various different environment configurations in Appendix E. Therefore, we conclude, that performance of the QRL and classical RL models is on par for the BeamManagement6G environment.
These closely match the ground-truth intensity distribution (yellow) obtained by brute force. Moreover, all approaches clearly improve upon a random strategy. Note that the policies of the trained models correspond to non-trivial behavior patterns. We stress that simply selecting the closest antenna is sub-optimal as highlighted by the complex spatial intensity patterns shown in Figure 12 in the appendix. Moreover, as highlighted in Figure 5, the position of antenna and the mobile phone are unknown to the agent and only learned implicitly.
@@ 7 Discussion of Quantum Advantage
In this section we discuss the relation of our benchmarking results of the previous section to potential quantum advantage. Let us define heuristic quantum advantage as the consistent and statistically significant (Definition 4.1) outperformance (Definition 4.2) of a quantum algorithm over the best (known) classical algorithm for a given task. Do our findings show signatures of heuristic quantum advantage? Not quite, because despite extensive hyperparameter optimization, we cannot guarantee that we compared the quantum model to the optimal classical one.

### v2

@@ 1 Introduction
In the search for heuristic quantum advantage in QRL, we encounter issues similar to classical RL: sensitivity to hyperparameter choices, various randomness sources, and even random seed and codebase (Henderson et al., 2018; Jordan et al., 2024). QRL faces similar, if not more severe, reproducibility issues than traditional RL (Bowles et al., 2024; Franz et al., 2023) due to additional randomness. Therefore, any claim of QRL’s superiority over classical RL should be taken with great care.
@@
In addition to common sources of randomness known in RL (Henderson et al., 2018), such as random weight initialization, randomness in the environment, etc., QRL is subject to additional sources of randomness such as shot noise or hardware imperfections due to current limitations. Therefore, performance comparisons need to be based on a sound statistical evaluation. However, most studies perform only a small, potentially insufficient number of training runs for robust inferences under stochasticity. As a result, statements are potentially misleading or insignificant. Figure 2 exemplifies the comparison of two RL algorithms w.r.t some threshold on the evaluated return. It illustrates common flaws prevalent in the QRL literature, e.g., averaging learning curves over only 5 seeds, inconsistency in statistical ranges, etc. Figure 2 indicates that algorithm 1 is more sample-efficient. However, the interquartile ranges (shaded areas) estimated with a much larger sample size of $100$ runs rather support the opposite statement.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work. Section 3 provides background on (Quantum) RL. Next, Section 4 introduces our statistical estimator. Section 5 describes the experimental setup and introduces a novel environment with flexible complexity. Section 6 evaluates QRL using our statistical estimator. We discuss the implications on quantum advantage in Section 7.
@@ 2 Related Work
For classical deep RL, in general, no rigorous analysis of algorithms is possible, necessitating a computational approach to algorithmic comparison and representative benchmarking environments (Brockman et al., 2016; Todorov et al., 2012; Tassa et al., 2018a) for mature implementations of widely-used deep RL algorithms (Hill et al., 2018; Duan et al., 2016; Fujimoto et al., 2019; Wang et al., 2019b). However, benchmarking results are heavily influenced by (sometimes seemingly trivial) implementation details (Engstrom et al., 2019; Andrychowicz et al., 2021; Huang et al., 2022), as well as randomness in environment transitions and network initialization (Henderson et al., 2018). Research has focused on determining the necessary number of seeds for robust comparison (Agarwal et al., 2021; Colas et al., 2018) and on adopting statistical analysis over simple point estimates (Agarwal et al., 2021; Colas et al., 2018; Patterson et al., 2024; Huang et al., 2024) (e.g., reporting the interquartile mean instead of the average performance). However, benchmarking algorithms with increased number of seeds (Laskin et al., 2019; Gorsane et al., 2022; Bettini et al., 2024) can quickly become a bottleneck (Jordan et al., 2024), especially for computationally-intense algorithms, like model-based (Wang et al., 2019b) or safe RL (Zhao et al., 2024). In addition, (Jordan et al., 2020) emphasizes that standard evaluations often overlook the difficulty of hyperparameter tuning, which can obscure the true usability of an algorithm in real-world applications.
@@
QRL can solve artificial tasks (without practical relevance) with exponentially smaller sample complexity compared to any known classical algorithm. Recently, (Jerbi et al., 2021; Liu et al., 2021) constructed artificial problems which are widely believed to be classically intractable. When the type of the algorithm is fixed to e.g. policy iteration, quantum versions exist with polynomially reduced sample complexity (Wang et al., 2021; Ganguly et al., 2023; Zhong et al., 2024; Wiedemann et al., 2023). However, these algorithms require resources that far exceed the capabilities of current quantum hardware. An alternative line of research (Chen et al., 2020; Skolik et al., 2022; Lockwood & Si, 2020), more aligned with the limitations of current hardware, substitutes the classical neural network used as a function approximator for policy and value function in classical RL algorithms by a variational quantum circuit VQC (Bharti et al., 2022). Here, Jerbi et al. (2021) construct artificial problems based on the same VQC architecture later used by the learner. The superiority found empirically for the QRL algorithm can then be attributed to the inductive bias introduced in the problem. Several studies have compared classical RL and QRL on toy problems inspired by real-world tasks. Some of these studies (Chen et al., 2020; Drǎgan et al., 2024; Reers, 2023; Hohenfeld et al., 2024; Eisenmann et al., 2024) have documented empirical superiority of QRL on metrics closely related to sample complexity. This work proposes a robust methodology to compare these types of algorithms.
@@
We compare two function approximation approaches in these algorithms: (i) the standard version with a classical DNN and (ii) the quantum version, which has a variational quantum circuit between two small classical fully-connected layers. This structure is called hybrid classical-quantum, as depicted in Figure 3. As discussed in Section D.2, the classical layers have few trainable parameters, and small fully classical DNNs are not competitive, so the hybrid models’ performance originates from the quantum sub-module. The VQC acts like a quantum counterpart to a classical DNN, with trainable weights parameterizing unitary operations, usually single-qubit rotations (Bharti et al., 2022). The quantum state is measured to approximate complex functions. Past research suggests VQCs may have advantages over DNNs, including better accuracy for certain tasks (Liu et al., 2021) and smaller model size (Chen et al., 2020). More details on VQCs and the ansatz used here are shown in Section D.1. Such models have been used in value-based (Chen et al., 2020; Skolik et al., 2022) and policy-based (Jerbi et al., 2021; Meyer et al., 2023a) RL routines.
@@ 5 Experimental Setup
As QRL algorithms exhibit some robustness against hardware noise (Skolik et al., 2023) (which is expected to further decrease in the future (Kim et al., 2023; Acharya et al., 2025)) we decided to perform all experiments in a noise-free simulation. We reduced the impact of parameter initialization by considering $100$ random seeds per environment and model configuration. We initialized the classical neural networks using He initialization and the VQC parameters uniformly at random within $\left[0,2\pi\right]$. We tuned the hyperparameters, see Appendix C, before running the actual experiments. The configurations of the models are reported in terms of three metrics: (i) model width, i.e., the number of neurons in hidden layers for classical DNNs, and in case of quantum architectures the number of qubits; (ii) model depth, i.e., the number of hidden layers for classical networks, and the number of layers of the quantum models; and (iii) model complexity, parameterized by the number of trainable parameters, which will be justified further below. The reported configurations were chosen by an extensive ablation study, see Appendix D.
@@
We specifically employed this environment because it has a small state and action space, yet exhibits complex dynamics, enabling meaningful analysis without overwhelming encoding complexity. Moreover, its state space is mostly continuous, reflecting realistic energy variations and beam selections as continuous angles. Finally, there is ongoing debate on suitable methods for beam management, with RL identified as a strong contender (Maggi et al., 2024; Voigt et al., 2025), making this environment an excellent choice for benchmarking classical RL and QRL in a setting inspired by practical 6G challenges. Nonetheless, we highlight that this environment is merely one example to empirically demonstrate our methodology, which can also be applied to other, entirely different tasks.
@@ 6 Experiments
We evaluate different (Q)RL algorithms on our novel BeamManagement6G environment using our proposed statistical sample complexity estimator in Section 6.1. Section 6.2 studies the relationship between sample complexity and (RL) model complexity. In Section 6.3 we compare the performance of the trained policies.
@@ 6.2 Scaling with Model Complexity
We now explore the relationship between sample efficiency and the complexity of the RL and QRL model, respectively. There exist different measures for model complexity both in the classical (Hu et al., 2021) and quantum (Abbas et al., 2021) domain. In this work we define model complexity based on the number of trainable parameters. The parameter count defines a sequence of models in model space (assuming additional hyperparamter search for given parameter count). This sequence can then be used to identify potential trends, for example in the behavior of sample complexity as we scale to more powerful quantum models.
@@
At this point, two questions arise which are addressed in the following: (i) Are models with lower complexity but similar performance generally preferrable, i.e., how should sample and model complexity be balanced? (ii) Is outperfomance of the quantum model achievable by further increasing qubit numbers? The first question (i) cannot be answered in general as it depends on concrete practical considerations and objectives. If the primary goal is to minimize the number of agent-environment interactions, regardless of training and inference costs, the classical models are superior. However, when memory requirements are considered, the reduced parameter count of the quantum-enhanced model might be advantageous. Assuming scalability of this approach, also runtime improvements are conceivable. The second question (ii) cannot be answered conclusively at the current stage. The saturation behavior in Figure 7 suggests that further scaling the quantum model could lead to comparable or superior performance relative to the largest classical model, which, however, can only be substantiated by more experimental evaluations. We stress that a naive extrapolation to higher qubit numbers might be limited in practice by trainability issues (Larocca et al., 2024). Moreover, it is computationally prohibitive to generate simulation results beyond $14$ qubits due to the substantial overhead required for statistically robust results as summarized in Table 3 in Section D.2.
@@ 6.3 Performance of Trained Policies
As a consequence of different trajectories, the maximum achievable intensity varies. Both RL strategies produce similar-looking histograms of the received intensities. The reported mean values of the intensity of the quantum model (green) and the classical model (blue) do not indicate significantly different behavior by visual inspection. Furthermore, we observed comparable results for various different environment configurations in Appendix E. Therefore, we conclude that the performance of the QRL and classical RL models is on par for the BeamManagement6G environment.
These closely match the ground-truth intensity distribution (yellow) obtained by brute force. Moreover, all approaches clearly improve upon a random strategy. Note that the policies of the trained models correspond to non-trivial behavior patterns. We stress that simply selecting the closest antenna is sub-optimal, as highlighted by the complex spatial intensity patterns shown in Figure 13 in the appendix. Moreover, as highlighted in Figure 5, the position of antenna and the mobile phone are unknown to the agent and only learned implicitly.
6.4 Standard CartPole Benchmark
(a) Surface plot of sample complexity over various $\epsilon$-$\delta$ configurations. The classical model contains one hidden layer of width $16$, the quantum model operates on $4$ qubits.
(b) A cross-cut of sample complexity at a relative error of $\varepsilon=0.05$. The error bars denote the 5th and 95th percentiles.
Figure 9: Comparing empirical sample complexities $\hat{S}$ of vanilla policy gradients and a quantum version of the algorithm on the CartPole-v1 environment. Both classical and quantum models were optimized over a wide range of hyperparameters and model sizes. The estimates are based on $100$ training runs.
To further test our benchmarking scheme, we extend our analysis to the widely used CartPole-v1 environment. This benchmark regularly appears in QRL studies  (Lockwood & Si, 2020; Skolik et al., 2022), often accompanied by claims that quantum models outperform classical approaches with fewer number of trainable parameters. However, these statements rely on a small number of training runs and are based on simple visual comparisons. In contrast, here we apply our statistical sample complexity estimator to CartPole-v1 with the results shown in Figure 9.
We employed a vanilla policy gradient algorithm and performed hyperparameter search over a similar range as reported in Section C.1. While generally we do not expect to find globally optimal settings (Jordan et al., 2020), more advanced training methods such as natural gradients (Meyer et al., 2023b) could potentially further improve performance. Nevertheless, the chosen configurations allows us to compare classical and quantum function approximators with the number of trainable parameters ranging from about $30$ to $17,000$. The best quantum setup found, a single-layer model on only four qubits (Meyer et al., 2023a), performs competitively to the best classical model across nearly all $\epsilon$-$\delta$ configurations. The cross-cut, Figure 9(b), of Figure 9(a) at an evaluated threshold of $\varepsilon=0.05$ shows that the performance gap may not always be statistically significant due to overlapping confidence bounds. This again underscores the importance of sufficient number of runs.
We stress that these results should not be interpreted as evidence for quantum advantage. The circuit sizes are very small, allowing straightforward classical simulation, and scaling CartPole-v1 in a way that requires larger quantum models is notoriously difficult. These limitations motivate our focus on BeamManagement6G, whose complexity can be flexibly scaled.
@@ 7 Discussion of Quantum Advantage
In this section, we discuss the relation of our benchmarking results of the previous section to potential quantum advantage. Let us define heuristic quantum advantage as the consistent and statistically significant (Definition 4.1) outperformance (Definition 4.2) of a quantum algorithm over the best (known) classical algorithm for a given task. Do our findings show signatures of heuristic quantum advantage? Not quite, because despite extensive hyperparameter optimization, we cannot guarantee that we compared the quantum model to the optimal classical one.
@@ Acknowledgements
Funding: This research was conducted within the Bench-QC project, a lighthouse project of the Munich Quantum Valley initiative, which is supported by the Bavarian state with funds from the Hightech Agenda Bayern Plus.
Impact Statement
This paper aims to advance the field of Quantum Reinforcement Learning by introducing a statistical benchmarking routine for rigorous evaluation. Our work addresses the need for substantiated claims of quantum advantage, promoting a standard of thorough assessment. We anticipate that this contribution will enhance the quality and reliability of future research in Quantum Reinforcement Learning, supporting the development of robust and verifiable advancements in the field.
@@
Statement of Independent Work
We acknowledge the existence of a research paper with the same title as ours (Kruse et al., 2025). We wish to clarify that both works were conducted independently and without knowledge of each other.